A group of soldiers is sent to the jungle on an expedition for 60 days. They carry food enough for 60 days at 1 2/3 lbs for each person a day. After 15 days, 10 of the men die from malaria, and the rest of the men have a daily allowance of 1 4/5 lbs. per person for the balance of the trip. Find the original number of men.
Answer by htmentor(1343)

Let n = the original number of men
The initial amount of food for the 60 day expedition = 5/3 lb/man/day*60*n = 100n
After 15 days n men have consumed 5/3 lb/day/man*15*n = 25n lbs of food
The food remaining is 75n and this is enough for n-10 men to have 1 4/5 lb per day for the 45 days left.
Therefore 9/5 = 75n/45/(n-10)
Solve for n:
9/5(n-10) = (15/9)n
n - 10 = 75/81n
(6/81)n = 10
n = 810/6 = 135
So the original number of men = 135